# GEDAN: Learning the Edit Costs for Graph Edit Distance

### v1

@@ Appendix A Details on the formulation of GEDAN
$\begin{aligned} \operatorname{S-GEDAN}(G,G^{\prime})=\mathbf{P}\odot\left[\lambda_{1}\left((\sum_{k=0}^{K}\mathbf{M}_{k})\cdot\frac{1}{\mathcal{T}}+\mathbf{B}+\mathbf{D}\right)+\lambda_{2}\left((\sum_{k=0}^{K}\left(\mathbf{M}_{k}\odot\mathbf{C}_{k}\cdot\beta_{k}\right))\cdot\frac{1}{\beta_{K}}\right)\right]+\lambda_{\mathcal{T}}\cdot log(\mathcal{T}).\end{aligned}$
where $\lambda_{\mathcal{T}}$ modulates the effect of the logarithm of the temperature $\mathcal{T}$.
A.1 Layer Influence and Inference Dynamics
@@
To isolate the effect of architectural depth, we conduct a first experiment using U-GEDAN in a purely inference setting, without any prior training. As shown in Table 3, increasing $K$ leads to both a higher parameter and consistent improvements in rank-based metrics $\tau$ and $\rho$. This suggests enhanced alignment with the underlying graph structure. However, this increase in structural coherence does not directly translate to better RMSE performance. We hypothesize that this is due to the nature of the injective $A^{k}$ layers, which are implemented as non-linear GIN layers. While such $K$ layers improve relational consistency, they do not necessarily minimize prediction error in an untrained setting.
Table 3: U-GEDAN inference of the GED approximation with $K$-layers on Configuration 1.
LAYERS AIDS Morris et al. [2020] MUTAG Morris et al. [2020] PTC MR Morris et al. [2020] K Avg. Param. $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ 1 6K $\mathbf{6.71\pm 0.07}$ $0.384\pm 0.07$ $0.600\pm 0.07$ $11.52\pm 0.74$ $0.602\pm 0.09$ $0.844\pm 0.05$ $\mathbf{4.92\pm 0.52}$ $0.585\pm 0.08$ $0.785\pm 0.06$ 2 8K $6.79\pm 0.13$ $0.393\pm 0.07$ $0.609\pm 0.07$ $11.77\pm 0.80$ $0.605\pm 0.09$ $0.848\pm 0.04$ $4.96\pm 0.52$ $0.598\pm 0.08$ $0.795\pm 0.06$ 4 12K $6.80\pm 0.11$ $0.404\pm 0.07$ $0.609\pm 0.07$ $11.40\pm 0.93$ $0.612\pm 0.09$ $0.855\pm 0.04$ $5.26\pm 0.57$ $0.605\pm 0.08$ $0.801\pm 0.06$ 8 19K $7.20\pm 0.23$ $0.412\pm 0.07$ $0.630\pm 0.07$ $10.60\pm 0.89$ $0.618\pm 0.08$ $0.862\pm 0.04$ $5.70\pm 0.56$ $0.614\pm 0.07$ $0.809\pm 0.06$ 16 35K $7.48\pm 0.14$ $0.422\pm 0.07$ $0.638\pm 0.07$ $10.15\pm 0.93$ $0.618\pm 0.08$ $0.868\pm 0.03$ $6.17\pm 0.60$ $0.621\pm 0.08$ $0.814\pm 0.06$ 32 66K $7.91\pm 0.21$ $\mathbf{0.430\pm 0.07}$ $\mathbf{0.647\pm 0.07}$ $\mathbf{9.29\pm 1.12}$ $\mathbf{0.615\pm 0.08}$ $\mathbf{0.871\pm 0.03}$ $6.64\pm 0.63$ $\mathbf{0.633\pm 0.08}$ $\mathbf{0.823\pm 0.06}$
@@
Figure 3: The training curves of U-GEDAN on the test-set of RMSE, $\tau$ and $\rho$ respectively.
An additional analysis can be performed by observing the scatter plot between the ground-truth GED and the predicted values across training epochs. Over the 25 training epochs, we observe the evolution in the model’s internal representations, which directly affects its predictions. In Fig. 4, we report the scatter plots at three stages of training: (a) epoch 1, (b) epoch 13, and (c) epoch 25. Each point represents a graph pair. The red arrow highlights a specific graph pair whose prediction significantly improves over time, moving closer to the identity line.
Figure 4: Scatter plots showing the GED predicted by U-GEDAN and ground-truth GED values at different training epochs: (a) epoch 1, (b) epoch 13, and (c) epoch 25.
@@
This improvement is achieved in an unsupervised way, without direct supervision on the GED values, showing that U-GEDAN can internally structure representations that aligns with true graph distances.
Appendix B Details of the GED Approximation
In this section, we report the results obtained for each configuration used in our GED approximation experiments. We selected three benchmark datasets Morris et al. [2020]: AIDS, MUTAG, and PTC MR (see Table 4).
@@ Appendix B Details of the GED Approximation
AIDS Morris et al. [2020] MUTAG Morris et al. [2020] PTC MR Morris et al. [2020] N. graph pairs Avg. nodes Avg. edges N. graph pairs Avg. nodes Avg. edges N. graph pairs Avg. nodes Avg. edges 190096 8.08 15.39 5776 13.29 28.08 16900 6.95 13.00
Furthermore, this setting allows us to build upon the recent analysis by Yang et al. [2024], which highlights the potential to achieve high-quality GED approximations even in low-label regimes. In particular, we use the MUTAG dataset to explore how models behave under data scarcity conditions. Specifically, the number of graph pairs used in our experiments is as follows: MUTAG contains 5576 pairs, PTC MR contains 16900 pairs, and AIDS contains 190096 pairs.
@@
B.1 Results For Each Configuration
In Tab. 6, 7, 8, 9, 10, we present the results obtained from each model on the three datasets. All results are obtained by cross-validation with 5-fold K-fold scheme, using the 10% of training set as validation set for model selection. For each fold, we select the model with the best performance on the validation set and then use it for inference on the test set.
A comparative analysis of the results across configurations confirms the absence of a universally superior model, consistent with the no-free-lunch theorem Wolpert and Macready [1997].

### v2

@@ Appendix B Formal Structures of the Matrices
$\begin{aligned} \operatorname{GEDAN}(G,G^{\prime})=\mathbf{P}\odot\left[\lambda\left((\sum_{k=0}^{K}\mathbf{M}_{k})\cdot\frac{1}{\mathcal{T}}+\mathbf{B}+\mathbf{D}\right)+(1-\lambda)\left((\sum_{k=0}^{K}\left(\mathbf{M}_{k}\odot\mathbf{C}_{k}\cdot\beta_{k}\right))\cdot\frac{1}{\beta_{K}}\right)\right]+\lambda_{\mathcal{T}}\cdot log(\mathcal{T}),\end{aligned}$
where $\lambda_{\mathcal{T}}$ modulates the effect of the logarithm of the temperature $\mathcal{T}$. The temperature regulates the “softness” of the matching matrices: higher values introduce greater uncertainty into the optimal matching, favoring exploration during optimization; conversely, lower values make the matrices more rigid, accelerating convergence but increasing the risk of converging towards local minima and suboptimal matches. The parameter $\lambda_{\mathcal{T}}$ helps stabilize optimization as a function of temperature.
Appendix C Model Details
Figure 4: The architectural scheme of GEDAN
As for the functions $f^{(k)}\rightarrow\mathbf{C}_{k}$, we concatenate the representations of the two nodes involved and provide the result as input to an MLP. This MLP employs a ReLU as an intermediate activation function and applies a softplus function with parameter $\beta=5$ on the final output, as described in the main body of the text.
Training GEDAN requires substantial computational resources. We employ a NVIDIA A100 GPUs for GED approximation, while experiments on learning edit costs are conducted on a cluster with two NVIDIA H100 GPUs to optimize training time and cost. The implementation is based on PyTorch (v2.2.0) together with the PyTorch Geometric library (v2.4.0), with hardware acceleration provided by CUDA (v12.1). The source code is publicly available at https://github.com/MonsieurSolver/GEDAN.
C.1 Illustrative Graph Matching Example with $M_{k}$ Penalties
In Fig. 1 we illustrate how our model evaluates the matching between two graphs. To provide a concrete example, we consider a real case from the PTC MR dataset (Morris et al., 2020), where two structurally similar graphs are compared. This example highlights how the additive penalty improves the quality of the matching.
Fig. 5 reports this real-world instance. For clarity of visualization, we restrict the matching matrix to the first eight nodes. In this setting, nodes 0 and 5 of graph $G$ can potentially match with nodes 1 and 4 of graph $G^{\prime}$. Importantly, we stress that the analysis here is based solely on the matrices $\mathbf{M}_{k}$, without directly considering node degrees.
Figure 5: Matching example on PTC MR using $\mathbf{M}_{k}$ matrices.
At layer $\mathbf{M}_{0}$, which relies exclusively on node labels, we observe that nodes 0 and 5 of $G$ perfectly match with nodes 1 and 4 of $G^{\prime}$, respectively. Moving to $\mathbf{M}_{1}$, where neighborhood information is incorporated, clear differences emerge. Specifically, node 0 in $G$ has exactly one pink neighbor, while node 5 has two. A similar configuration is present in $G^{\prime}$, with node 1 having one pink neighbor and node 4 having two. Consequently, in the matching matrix, entry $\mathbf{M}_{1_{0,1}}$ exhibits a bright yellow color (distance 0), while $\mathbf{M}_{1_{0,4}}$ appears green (higher distance). Conversely, node 5 in $G$ yields a bright yellow value at $\mathbf{M}_{1_{5,4}}$, reflecting a nearly identical representation.
Proceeding to $\mathbf{M}_{2}$, additional refinements occur: similarity scores change for most nodes, while only the pair of nodes labeled 2 in both graphs retain a strong correspondence. Finally, when aggregating $\mathbf{M}_{T}=\sum_{i=0}\mathbf{M}_{i}$, the result clearly emphasizes the bright yellow values at $\mathbf{M}_{T_{0,1}}$ and $\mathbf{M}_{T_{5,4}}$, thereby reinforcing the correct matchings.
C.2 Impact of the Temperature $\mathcal{T}$
As shown in Formula 18, the approximation can be extended with a temperature parameter $\mathcal{T}$, which regulates the learning of the optimal matching. This is particularly useful since the parameter $K$ cannot always cover the maximum depth of the graphs. Allowing the model an initial exploration by softening the distance matrices enables the selection of non-ideal node pairs at early stages. Although this may appear counterintuitive, GIN networks update their representations differently under such conditions, and non-optimal initial matchings can lead to better solutions, precisely because the full graph structure is not observable.
To illustrate this effect, we consider graphs from the ENZYMES dataset (Morris et al., 2020) with up to 64 nodes. We employ a $128\times 128$ Gumbel-Sinkhorn module and restrict the number of layers to $k=3$. In this setting, a significant portion of the graph structure remains unobserved. Different values of $\mathcal{T}$ soften the matching matrix to varying degrees, thus promoting different exploration strategies and optimization trajectories. Performance is evaluated against the lower bound of the GED, defined as the minimum number of operations that must occur when transforming one graph into another, i.e., the absolute difference in the number of nodes and edges.
Figure 6: Effect of the temperature $\mathcal{T}$ on graph matching, showing how softening the distance matrix enables different exploration strategies and optimization outcomes.
C.3 The Loss to Approximate the Graph Edit Distance
As described in the main text, $\text{GEDAN}(\lambda=1)$ is trained in an unsupervised manner to approximate the Graph Edit Distance (GED) by optimizing its internal node representations without relying on ground-truth distances. The associated loss function is defined as
$\omega^{*}=\arg\min_{\omega}\sum f_{\omega}(G,G^{\prime}),$ (20)
where $\omega$ denotes the parameters of the GINs generating node representations $h^{(k)}(v)$ and $h^{(k)}(v^{\prime})$ for nodes $v\in G$ and $v^{\prime}\in G^{\prime}$, respectively. GEDAN ($f_{\omega}$) estimates the GED by leveraging the fact that it is zero if and only if two graphs are identical (see Eq. 10). This property provides an intrinsic learning signal, enabling unsupervised training.
This signal underpins the stability of learning: for two graphs $G$ and $G^{\prime}$ that are identical up to the $k$-WL test, there exists a node-wise matching such that corresponding representations at layer $k$ are identical, i.e., their pairwise distances are zero. The Gumbel-Sinkhorn module naturally selects these pairs as having minimal distance, preventing the model from altering them.
The multi-scale mechanism further reinforces meaningful alignment by penalizing differences between node representations: distances vary between 0 (identical) and 1 (maximally different). Nodes not belonging to the same graph incur higher assignment costs, while nodes that remain similar across scales up to layer $k$ are assigned lower costs and thus preferentially matched. This effectively encourages alignment of embeddings that are intrinsically similar, as measured by multi-scale distances.
To ensure stability and avoid trivial solutions, we impose two constraints on the learned representations. First, we use spherical embeddings to prevent the model from minimizing pairwise distances simply by shrinking norms. Second, we pre-train the assignment matrix $\mathbf{P}$, since training it from scratch can introduce gradient noise or allow trivial shortcuts. Pre-training ensures that the network has already learned a discrete matching strategy; freezing $\mathbf{P}$ during subsequent training forces the network to minimize the loss solely through the embeddings. Moreover, to prevent other shortcuts, we do not optimize the Gumbel-Sinkhorn network nor the deletion/insertion costs $a^{\ominus},a^{\oplus},b^{\ominus},b^{\oplus}$, effectively setting $\lambda=1$ and excluding the use of matrices $\mathbf{C}_{k}$. Without these constraints, the network could trivially minimize the loss by adjusting these components, bypassing meaningful representation learning.
In summary, this framework encourages alignment of similar nodes by minimizing the total assignment cost through their embeddings. The matching mechanism itself remains fixed; embeddings “self-organize” under gradient descent to reduce the total cost. The loss attains a guaranteed minimum of zero if and only if the two graphs are identical, or equivalently, indistinguishable under the $k$-Weisfeiler-Lehman test.
C.4 Gumbel-Sinkhorn
To obtain a good $\mathbf{P}$, we modify the version of the original Gumbel-Sinkhorn network (Mena et al., 2018; Shah et al., 2024), using the Truncated Gumbel (Neamah & Qasim, 2021) to reduce the gradient variance and add stability. We then sample with the Gumbel-Max Trick (Huijben et al., 2023) while maintaining differentiability. Finally we stabilize the normalization of Sinkhorn (Sinkhorn, 1964) with an early stop, improving efficiency and limiting gradient vanishing (LeCun et al., 1998).
The Gumbel-Sinkhorn network has been trained for 200 epochs on matrices of sizes $32\times 32$, $64\times 64$, $128\times 128$ and $256\times 256$. The matrices used for training are randomly generated, with 250,000 samples for each of the three dimensions. For each matrix, Linear Sum Assignment (LSA) is calculated and used as a reference label. Optimization is performed by minimizing the Huber Loss, with a learning rate of 0.001 and using the Adam optimizer. Training also includes a curriculum learning phase, during which many matrices with zero values on the diagonal are initially proposed, a strategy that facilitated model convergence. In addition, an early stopping mechanism is applied on the Sinkhorn process, with a limit of 50 iterations, where the stopping condition is determined by a difference between consecutive iterations of less than 1e-4.
To test the performance of the Gumbel-Sinkhorn network, 10000 new random matrices are generated, for which the LSA varies between 0.001 and 199.750, with an average of 77.375. The overall performance of the model calculated on the $64\times 64$ matrices using high temperatures has a $\tau$ value of 0.991, a $R^{2}$ coefficient of 0.993 and a Root Mean Square Error (RMSE) of 3.940. These configurations are kept stable in the pre-trained versions in GEDAN.
To assess the importance of pre-training (PT), we reproduce the results of Table 10 using a Gumbel-Sinkhorn module without pre-training (No-PT). In this setting, the network bears a heavier optimization load: although the model attempts to minimize the matching while jointly training the module, the dual optimization produces noisier and less accurate gradients. This effect is particularly critical in the unsupervised setting, where the model must self-organize; noisy gradients can destabilize the process and lead to highly suboptimal solutions. In contrast, under supervised training the impact is less severe, as shown in GraphEdXXOR (Jain et al., 2024), where the module is not pre-trained but remains viable thanks to the stable guidance provided by labels.
Table 3: Ablation study on the effect of pre-training (PT) the Gumbel-Sinkhorn module versus no pre-training (No-PT) on Configuration 1. Best results for each metric and dataset are highlighted in boldface.
MODELS AIDS [38] MUTAG [38] PTC MR [38] $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ PT GEDAN($\lambda=0$) $\mathbf{2.82\pm 0.19}$ $\mathbf{0.484\pm 0.05}$ $\mathbf{0.694\pm 0.05}$ $\mathbf{5.20\pm 0.15}$ $\mathbf{0.650\pm 0.04}$ $\mathbf{0.887\pm 0.01}$ $\mathbf{3.18\pm 0.65}$ $\mathbf{0.655\pm 0.08}$ $\mathbf{0.826\pm 0.07}$ GEDAN($\lambda=1$) $\mathbf{3.54\pm 0.14}$ $0.437\pm 0.06$ $0.652\pm 0.06$ $\mathbf{13.30\pm 0.85}$ $\mathbf{0.611\pm 0.09}$ $\mathbf{0.852\pm 0.04}$ $\mathbf{3.51\pm 0.31}$ $\mathbf{0.649\pm 0.07}$ $\mathbf{0.823\pm 0.05}$ No-PT GEDAN($\lambda=0$) $6.973\pm 3.89$ $0.193\pm 0.10$ $0.254\pm 0.14$ $916.82\pm 8.83$ $0.603\pm 0.05$ $0.723\pm 0.05$ $5.52\pm 1.75$ $0.445\pm 0.09$ $0.591\pm 0.08$ GEDAN($\lambda=1$) $2509.84\pm 18.5$ $0.074\pm 0.04$ $0.101\pm 0.07$ $3224.39\pm 33.8$ $0.596\pm 0.05$ $0.722\pm 0.05$ $2335.98\pm 121.9$ $0.396\pm 0.07$ $0.584\pm 0.08$
The computational complexity of the Gumbel–Sinkhorn module is $O(Ln^{2})$, where $L$ denotes the number of Sinkhorn normalization iterations. This is lower than the complexity of the Kuhn–Munkres algorithm ($O(n^{3})$(Kuhn, 1955). In the case of our model (GEDAN), choosing the number of iterations such that $L\ll n$ keeps the computational cost manageable even for large matrices. Recent works, such as Tang et al. (2024), further demonstrate that this complexity can be reduced to $\approx O(n^{2})$, providing strategies that are particularly useful for improving the efficiency of our approach.
C.5 Influence of Depth and Inference Dynamics
@@
To isolate the effect of architectural depth, we conduct a first experiment using $\text{GEDAN}(\lambda=1)$ in a purely inference setting, without any prior training. As shown in Table 4, increasing $K$ leads to both a higher parameter and consistent improvements in rank-based metrics $\tau$ and $\rho$. This suggests enhanced alignment with the underlying graph structure. However, this increase in structural coherence does not directly translate to better RMSE performance. We hypothesize that this is due to the nature of the injective $A^{k}$ layers, which are implemented as non-linear GIN layers. While such $K$ layers improve relational consistency, they do not necessarily minimize prediction error in an untrained setting.
Table 4: $\text{GEDAN}(\lambda=1)$ inference of the GED approximation with $K$-layers on Configuration 1.
LAYERS AIDS [38] MUTAG [38] PTC MR [38] K Avg. Param. $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ 1 6K $\mathbf{6.71\pm 0.07}$ $0.384\pm 0.07$ $0.600\pm 0.07$ $11.52\pm 0.74$ $0.602\pm 0.09$ $0.844\pm 0.05$ $\mathbf{4.92\pm 0.52}$ $0.585\pm 0.08$ $0.785\pm 0.06$ 2 8K $6.79\pm 0.13$ $0.393\pm 0.07$ $0.609\pm 0.07$ $11.77\pm 0.80$ $0.605\pm 0.09$ $0.848\pm 0.04$ $4.96\pm 0.52$ $0.598\pm 0.08$ $0.795\pm 0.06$ 4 12K $6.80\pm 0.11$ $0.404\pm 0.07$ $0.609\pm 0.07$ $11.40\pm 0.93$ $0.612\pm 0.09$ $0.855\pm 0.04$ $5.26\pm 0.57$ $0.605\pm 0.08$ $0.801\pm 0.06$ 8 19K $7.20\pm 0.23$ $0.412\pm 0.07$ $0.630\pm 0.07$ $10.60\pm 0.89$ $0.618\pm 0.08$ $0.862\pm 0.04$ $5.70\pm 0.56$ $0.614\pm 0.07$ $0.809\pm 0.06$ 16 35K $7.48\pm 0.14$ $0.422\pm 0.07$ $0.638\pm 0.07$ $10.15\pm 0.93$ $0.618\pm 0.08$ $0.868\pm 0.03$ $6.17\pm 0.60$ $0.621\pm 0.08$ $0.814\pm 0.06$ 32 66K $7.91\pm 0.21$ $\mathbf{0.430\pm 0.07}$ $\mathbf{0.647\pm 0.07}$ $\mathbf{9.29\pm 1.12}$ $\mathbf{0.615\pm 0.08}$ $\mathbf{0.871\pm 0.03}$ $6.64\pm 0.63$ $\mathbf{0.633\pm 0.08}$ $\mathbf{0.823\pm 0.06}$
@@
Figure 7: The training curves of $\text{GEDAN}(\lambda=1)$ on the test-set of RMSE, $\tau$ and $\rho$ respectively.
An additional analysis can be performed by observing the scatter plot between the ground-truth GED and the predicted values across training epochs. Over the 25 training epochs, we observe the evolution in the model’s internal representations, which directly affects its predictions.
In Fig. 8, we report the scatter plots at three stages of training: (a) epoch 1, (b) epoch 13, and (c) epoch 25. Each point represents a graph pair. The red arrow highlights a specific graph pair whose prediction significantly improves over time, moving closer to the identity line.
@@
Figure 8: Scatter plots showing the GED predicted by $\text{GEDAN}(\lambda=1)$ and ground-truth GED values at different training epochs: (a) epoch 1, (b) epoch 13, and (c) epoch 25.
This improvement is achieved in an unsupervised way, without direct supervision on the GED values, showing that $\text{GEDAN}(\lambda=1)$ can internally structure representations that aligns with true graph distances. Further examples are shown in Figures 121314 and 15.
Appendix D Details of the GED Approximation
In this section, we report the results obtained for each configuration used in our GED approximation experiments. We selected three benchmark datasets (Morris et al., 2020): AIDS, MUTAG, and PTC MR (see Table 5).
@@ Appendix D Details of the GED Approximation
AIDS MUTAG PTC MR N. graph pairs Avg. nodes Avg. edges N. graph pairs Avg. nodes Avg. edges N. graph pairs Avg. nodes Avg. edges 190,096 8.08 15.39 5,776 13.29 28.08 16,900 6.95 13.00
Furthermore, this setting allows us to build upon the recent analysis by Yang et al. (2024), which highlights the potential to achieve high-quality GED approximations even in low-label regimes. In particular, we use the MUTAG dataset to explore how models behave under data scarcity conditions. Specifically, the number of graph pairs used in our experiments is as follows: MUTAG contains 5,776 pairs, PTC MR contains 16,900 pairs, and AIDS contains 190,096 pairs.
D.1 The Setting Up of the Configurations
@@
For completeness, we report the statistics of GED values for each configuration and dataset considered. Recent work indicates that GED benchmarks may have some bias related to isomorphism, or that this bias may be easily introduced during the training phase. In our case, since the main goal is to compare our model with those in the literature, we acknowledge this potential criticality but leave it explicitly as a cue for future work. However, it should be emphasized that all datasets used are public and thus available for further analysis.
Table 7: Statistics of the AIDS dataset used in the GED approximation.
Dataset Configurations Mean Median Std Variance Min Max AIDS (Morris et al., 2020) 1 11.4 11.0 3.73 13.95 0.0 31.0 2 13.3 13.0 4.83 23.33 0.0 37.0 3 20.5 20.0 7.52 56.67 0.0 59.0 4 19.6 18.0 7.15 51.08 0.0 58.0 5 17.9 17.0 7.48 55.93 0.0 58.0
Figure 9: Statistics of the AIDS dataset.
Table 8: Statistics of the MUTAG dataset used in the GED approximation.
Dataset Configurations Mean Median Std Variance Min Max MUTAG (Morris et al., 2020) 1 30.9 31.0 8.90 79.24 0.0 51.0 2 15.9 16.0 4.47 20.01 0.0 26.0 3 19.3 20.0 5.33 28.39 0.0 32.0 4 22.3 22.0 6.65 44.22 0.0 40.0 5 20.0 20.0 5.74 32.97 0.0 37.0
Figure 10: Statistics of the MUTAG dataset.
Table 9: Statistics of the PTC MR dataset used in the GED approximation.
Dataset Configurations Mean Median Std Variance Min Max PTC MR (Morris et al., 2020) 1 13.0 13.0 5.29 28.00 0.0 30.0 2 15.3 15.0 6.61 43.80 0.0 38.0 3 21.3 21.0 9.26 85.69 0.0 53.0 4 19.6 18.0 9.92 98.55 0.0 60.0 5 19.5 18.0 10.0 100.9 0.0 60.0
Figure 11: Statistics of the PTC MR dataset.
D.2 Results For Each Configuration
In Tab. 10, 11, 12, 13, 14, we present the results obtained from each model on the three datasets. All results are obtained by cross-validation with 5-fold K-fold scheme, using the 10% of training set as validation set for model selection. For each fold, we select the model with the best performance on the validation set and then use it for inference on the test set.
We optimize each model by searching for configuration settings tailored to each dataset, varying both depth and number of parameters to balance performance quality and execution time. Although this optimization is not the primary focus of our work, it provides a useful estimate of how our approach diverges from prior literature on this specific task.
The only model not subject to parameter-level optimization is EUGENE (Bommakanti et al., 2024). This is because it does not offer control over parameters beyond cost settings, which we follow exactly as specified by the original authors. Another limitation of EUGENE (Bommakanti et al., 2024) is the inability to assign non-uniform costs to edges: the code defines a single edge cost, without differentiating between insertion and deletion, unlike nodes, for which separate costs can be specified for insertion, substitution, and deletion. To address this, we use an average cost, which prevents a fully comparable evaluation with the last two configurations.
Despite this, we observe that EUGENE (Bommakanti et al., 2024) remains, on average, less accurate than GEDAN even in configurations where cost settings are consistent. Furthermore, as shown later, EUGENE (Bommakanti et al., 2024) requires substantially more time for prediction. For these three reasons, summarized in Table 2, we prefer the GEDAN configuration with $\lambda=1$.
A comparative analysis across different configurations confirms the absence of a universally superior model, in agreement with the no-free-lunch theorem (Wolpert & Macready, 1997). The best performance is highlighted in boldface, while the second-best is underlined. For EUGENE (Bommakanti et al., 2024), we report the standard derivation, as we applied folding to accelerate the approximation using fewer graphs. In contrast, for the approximation proposed by Fischer et al. (2017), we approximated the entire dataset directly, since this method is efficient and, like EUGENE, does not require any training.